OpenStudy (anonymous):

If c is the measure of the hypotenuse and a and b are the measures of the legs, find the missing measure. Round the answer to the nearest tenth if necessary: a = 10, c = 21, b =?

OpenStudy (anonymous):

use Pythagorean theorem to figure this out ;D

OpenStudy (anonymous):

a^2 + b^2 = c^2 then solve for b^2 then b = ...